please help me solve this: Solve these problems by elimination
1)3x-y=26 x+y=10
2)135x+45y=-11 81x-75y=13
3)x+2y=3 2x+4y=6
4)3x-5y=50 2x+4y=-18

Answer by checkley77(12844)

1)
3x-y=26
x+y=10 add
-------
4x=36
x=36/4
x=9 ans.
3*9-y=26
27-y=26
-y=26-27
-y==1
y=1 ans.
---------------------------------------------
2)
135x+45y=-11 multiply this equation by 5
81x-75y=13 multiply this equation by 3 & add them
675x+225y=-55
243x-225y=39
-----------
918x=-16
x=-16/918
x=-8/459 ans.
81(-8/459)-75y=13
-648/459-75y=13
-75y=13+648/459
-75y=13+24/17
-75y=(13*17+24)/17
-75y=(221+24)/17
-75y=245/17
y=245/17*1/-75
y=-245/1275
y=-49/255 ans.
3)
x+2y=3 multiply by -2 & add
2x+4y=6
-2x-4y=-6
---------
0+0=0 these are identical equations with 2 unknowns. no unique solution.
-----------------------------------------------------------
4)
3x-5y=50 multiply by 4
2x+4y=-18 multiply by 5 & add
12x-20y=200
10x+20y=-90
-----------
22x=110
x=110/22
x=5 ans.
3*5-5y=50
15-5y=50
-5y=50-15
-5y=35
y=35/-5
y=-7 ans.